# Mil2: Efficient Cloth Simulation Using Non-distance Barriers and Subspace Reuse

### v3

@@ 3. Background
To make the exposition self-contained, we start with a brief review of projective dynamics and the distance-based interior point method for collision resolution. The reader can find more details from the relevant literature i.e. (Li et al., 2020; Lan et al., 2022b; Bouaziz et al., 2014).
@@
Here, $x$ is the unknown variable we would like to compute at the time step i.e., the position of all the cloth vertices. We also have:
(2) $z=x^{*}+h\dot{x}^{*}+h^{2}\mathsf{M}^{-1}f_{\text{ext}},$
as a known vector based on the previous displacement $x^{*}$, velocity $\dot{x}^{*}$, and an external force $f_{\text{ext}}$. $\mathsf{M}$ is the mass matrix, and $h$ is the time step size. The target function $E$ consists of the inertia momentum ($I$) offering a mass-weighted regularization over $x$, and the elasticity potential ($\Psi$) controlling the deformation of the cloth.
@@
The global stage is followed by local projections as a standard linear solve:
@@
Intuitively, the goal of Eq. (4) is to blend duplicated DOFs $y_{i}$ to produce a global solution of $x$ since $x_{i}$ could have multiple replicates if it is involved in several constraints. As a result, $w_{i}$ in Eq. (4) embodies the priority of a constraint – a bigger $w_{i}$ (relative to other constraints) means the global solve favors $x$ being closer to the corresponding $y_{i}$. In an extreme case when $w_{i}\rightarrow\infty$, $x_{i}\rightarrow y_{i}$ ignoring all the other constraints. One should not confuse $w_{i}$ with the stiffness of the constraint. High constraint stiffness produces a big internal force, which could overshoot and must be coupled with a line search for extra safeguards.
@@
Here, $\hat{d}$ is a user-provided tolerance of the collision resolution. $d_{i}(x)$ denotes the closest distance between the $i$-th pair of surface primitives, either a vertex-triangle pair or an edge-edge pair. $B(d_{i})$ diverges if $d_{i}<\hat{d}$ and approaches $\infty$ when $d_{i}\rightarrow 0$. Consequently, as long as we keep $d_{i}$ positive at the beginning of a time step (i.e., all primitives are separate), the existence of $B(d_{i})$ prevents any future inter-penetration with an increasingly stronger repulsion. Lan et al. (2022b) showed that IPC can also be integrated in LG iterations by setting $w_{i}$ of the collision constraint as $B(d_{i})$.
@@ 4. Non-distance Barrier
While IPC and its variations offer robust treatment for collisions, they are all based on $d_{i}$. Here, we name this family of barrier functions as the distance-based barrier or DBB. Updating DBB is expensive – one needs to perform a broad phrase collision culling to generate the list of vertex-triangle or edge-edge pairs, compute TOI for each pair, identify the smallest TOI that is between zero and one of all pairs, compute $d_{i}$ and eventually obtain the latest value of $\sum B(d_{i})$. Such a full CCD procedure is invoked frequently during LG iterations (see (Lan et al., 2022b)) and becomes the dominant computation along the pipeline.
@@
We argue that not all DBBs are indispensable, and most of them can be replaced with a more economic alternative model under the PD framework. The underlying philosophy of DBB is to offer a nonlinear penalty, which becomes stiffer when the collision constraint is about to be violated. It geometrically approximates the indicator functions ($\delta_{\mathcal{A}}(x)=0$, if $x\in\mathcal{A}$; $\delta_{\mathcal{A}}(x)=\infty$, if $x\notin\mathcal{A}$). However, a DBB itself is not physically accurate i.e., the gradient of DBB differs from the actual collision force unless $\kappa$ in Eq. (6) is sufficiently close to zero (the complementary slackness is fully satisfied). In practice, what we want is an increasingly strong repulsion to correct the constraint violation. Given the fact that one always needs multiple LG iterations to solve the barrier-in-the-loop PD problem, we define an exponential barrier (EXP) based on the how hard to resolve a penetrating primitive, and use EXP as $w_{i}$ of this constraint:
(7) $w_{i}=B_{i}^{EXP}=kA^{a},$
@@
where $a\in\mathbb{Z}^{+}$ represents how many consecutive LG iterations has the constraint been active. $k$ is an initial weight, and $A$ is the base of the growth rate. Eq. (7) suggests $B_{i}^{EXP}$ quickly approaches $\infty$ if a collision constraint $C_{i}$ remains unresolved over few LG iterations. When this is the case, the local target position of the constraint shall be satisfied with the highest priority to resolve the corresponding collision. $B_{i}$ in Eq. (7) no longer depends on the distance between primitives or any physically/geometrically meaningful measures. It becomes a self-adjusting variable as the optimization proceeds with minimum computation costs. As we elaborate in the next subsection, this feature delivers substantial convenience and speedup for cloth simulation.
@@ 4.1. Partial CCD
An immediate benefit of EXP is relieved CCD computation. It is easy to see from Eq. (7) that updating $B(x)$ only checks if a collision remains active after the previous LG iteration i.e, a true-or-false answer suffices, not the exact TOI. This feature allows a simplified approach for EXP CCD, which converts the cubic root-finding problem to multiple dot products, and we name this method partial CCD.
@@
As the cloth is less stretchable, we assume that edge lengths of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ are bounded by $L$. Under the finite velocity, it is reasonable to assume the distance between $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ is bounded above by $H^{1}$ during $t\in(0,1]$, and bounded below by $H^{0}$ at $t=0$. We can then obtain an upper bound of the norm of $\nabla_{\lambda}Q$:
@@
That said when $\rho_{\mathcal{S}}$ is smaller than $\rho^{*}$, there is at least one sample point sitting in a neighborhood of $Q<0$. Here, $\alpha<1$ is the hyperparameter used in line search filtering, which is typically set as $0.8$.
@@ Discussion
Sampling-based partial CCD can be considered an inexact approximate of exact CCD algorithms (Brochu et al., 2012), given the premise that EXP is in use. Eq. (15) provides a sufficient condition eliminating false negative collision detections. Nevertheless, false positive detections are still possible. Fortunately, the main objective of partial CCD is to estimate if a collision constraint remains alive. False positives do not affect the simulation quality or performance. Mil2 does not exclusively rely on partial CCD, and regular CCDs are still performed during the simulation (but at a much lower frequency). Therefore unless partial CCD is largely inaccurate, simulation performance and quality are not sensitive to partial CCD accuracy (e.g., see Fig. 10).
Eq. (15) appears sensitive to $H^{0}$, the cloest distance between $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ for $t\in(0,1]$. For instance, if a vertex is in the proximity of the triangle at $t=0$ and hits the triangle at $t=t^{*}$, $\rho(\lambda^{*})$ of the corresponding contact approaches zero i.e., the disc of $\Omega(\lambda^{*})$ shrinks to a point. This issue can be easily fixed by setting the projection of the vertex on the triangle as a sample point if they are close to each other. In some extreme cases, where the cloth is substantially stretched or accelerated, we increase the sample density by scaling the sample interval by the change ratio of $\rho^{*}$.
A quick benchmark well demonstrates the potential of such a sample-based strategy: solving a cubic equation on the GPU as in the original IPC implementation (Li et al., 2020) is slower than evaluating Eq. (9) on a RTX 3090 GPU by over $6,000$ times. For a high-resolution cloth model, we do not need a lot of sample points on a primitive, making partial CCD faster than regular CCD by orders.
@@ 5. Subspace Reuse
Exp enables a more affordable CCD processing, but it does not alter the fact that the optimization of Eq. (5) is highly nonlinear. Solving the system within a limited time budget remains a computational challenge for time-sensitive applications. Under the framework of PD, the bottleneck is the linear solve at the global stage i.e., Eq. (4). Commonly used strategies resort to iterative GPU solvers such as Jacobi (Wang, 2015) or Gauss-Seidel (Fratarcangeli et al., 2016) to solve Eq. (4) inexactly. Unfortunately, the presence of barriers makes those solvers less beneficial. Specifically, it is known that iterative solvers are effective in smoothing high-frequency errors but become cumbersome when dealing with low-frequency residuals. In contrast, low-frequency cloth deformation can be efficiently handled using subspace methods. A concrete example is shown in Figs. 2 and 3. To this end, we tackle Eq. (4) using both the reduced direct solver and fullspace iterative solver, aiming to reap the goods from both sides.
@@
(16) $\mathsf{U}^{\top}\mathsf{H}\left(X+\mathsf{U}q\right)=\mathsf{U}^{\top}b\;\Rightarrow\;\left(\mathsf{U}^{\top}\mathsf{H}\mathsf{U}\right)q=\Lambda q=\mathsf{U}^{\top}\left(b-\mathsf{H}X\right),$
@@
Figure 2. Bending beams w. and w/o subspace.  We simulate a collision-free scene where a cloth strip bends under gravity. The model consists of $30$K DOFs. The resulting deformation is of low-frequency, which is challenging for Jacobi-like methods. A subspace solve effectively resolves this issue: only $20$ A-Jacobi iterations are needed to fully converge the simulation, which otherwise takes over one thousand iterations. Because the bending stiffness is quite strong in this example, we need to assign a big SOR-like weight ($\omega=0.9$) to dampen each A-Jacobi iteration. Mil2 runs over $300$ FPS for this example, while PD-IPC (Lan et al., 2022b) is less than 0.5 FPS due to the large number of A-Jacobi iterations.
@@
Figure 3. Spectral distribution of residual errors (w/o collision).  We plot the distribution of residual error over the first 100 modal bases of the beam test (Fig. 2). As shown at the top, the dominant deformation is low-frequency, which is conveniently solved within the subspace (bottom). On the other hand, A-Jacobi iterations are not effective in dealing with low-frequency residuals. As shown in the middle, 100 A-Jacobi iterations barely lower the low-frequency errors, while the high-frequency strains are well relaxed.
@@
As $\mathsf{U}$ is eigenvectors, $\mathsf{U}^{\top}\mathsf{H}\mathsf{U}=\Lambda$ becomes a diagonal matrix of eigenvalues. If the constraint set stays unchanged during the simulation, solving Eq. (16) is highly efficient on the GPU – we only need one subspace projection for evaluating $\mathsf{U}^{\top}(b-\mathsf{H}X)$, and diagonalized subspace solve is neglectable. The resulting subspace displacement of $q$ is then converted to its fullspace counterpart as $u=\mathsf{U}q$. At this point, most low-frequency errors have been eliminated by the subspace solve, and $\widetilde{x}=X+\mathsf{U}q$ represents an ideal guess of Eq (4) for iterative solvers, which only has high-frequency errors. We then use aggregated-Jacobi (A-Jacobi) as in (Lan et al., 2022b) to solve for $\Delta x$ to relax the (high-frequency) residual of $\widetilde{x}$:
(17) $\mathsf{H}(\widetilde{x}+\Delta x)=b\;\Rightarrow\;\mathsf{H}\Delta x=b-\mathsf{H}\widetilde{x}=b-\mathsf{H}(X+\mathsf{U}q).$
@@
A representative experiment is reported in Fig. 2, where a cloth strip is attached to the wall and gets bent under gravity. When the low-frequency deformation is computed within the subspace ($r=30$), it only takes 20 A-Jacobi iterations to converge the simulation to the ground truth. General A-Jacobi iterations are not effective for such low-frequency deformations – we observe noticeable visual difference even after 600 iterations (the right-most beam). In this case, the bending stiffness of the cloth is relatively high, and the vanilla Jacobi or A-Jacobi do not even converge. We use a large SOR-like weight ($\omega=0.9$) to dampen each A-Jacobi update. The error distribution of this experiment is plotted in Fig. 3, which is consistent with our previous analysis. Collisions are ubiquitous in cloth simulation, and $\mathsf{H}$ constantly varies during the simulation. Building a new subspace at the simulation runtime does not sound practical for interactive simulations. Our key observation here is: while the global-stage matrix is altered by different collision constraints, the essential structure of its low-frequency subspace remains unchanged. It seems counterintuitive at first sight: how can old eigenvectors still be effective while the matrix is modified? This is because low-frequency modes built at the rest shape depict general and global cloth motions. Those motions are not strongly coupled with high-frequency deformations. For example, the distinct geometries of local wrinkles are influenced less by the overarching movements of the cloth and more by the specific configurations of the corresponding collisions. Interestingly, it is the low-frequency deformations that most significantly hinder the convergence of iterative solvers. By eliminating or even just reducing these low-frequency errors, we could significantly enhance the convergence of the subsequent Jacobi method.
@@
Figure 5. Subspace reuse.  A piece of tablecloth drops on a wooden Armadillo. There are 66K DOFs on the cloth, and over $30\%$ of them are involved in collision constraints.
@@
Figure 6. Convergence curves w. and w/o subspace (reuse).  We plot the convergence curves for experiments in Figs. 2 and 5. We can see subspace solve saves a large faction of A-Jacobi iterations even under intensive collisions. The computation cost of one subspace solve, on the other hand, is similar to performing one or two A-Jacobi iterations on average.

### v4

@@ 3. Background
To make the exposition self-contained, we start with a brief review of projective dynamics and the distance-based interior point method for collision resolution. The reader can find more details from the relevant literature e.g., (Li et al., 2020; Lan et al., 2022b; Bouaziz et al., 2014).
@@
Here, $x$ is the unknown variable we would like to compute at the next time step i.e., the position of all the cloth vertices. We also have:
(2) $z=x^{*}+h\dot{x}^{*}+h^{2}\mathsf{M}^{-1}f_{ext},$
as a known vector based on the previous position $x^{*}$, velocity $\dot{x}^{*}$, and an external force $f_{ext}$. $\mathsf{M}$ is the mass matrix, and $h$ is the time step size. The objective function $E$ consists of the inertia momentum ($I$) offering a mass-weighted regularization over $x$, and the elasticity potential ($\Psi$) controlling the deformation of the cloth.
@@
The global stage follows as a standard linear solve:
@@
Intuitively, the goal of Eq. (4) is to blend duplicated DOFs $y_{i}$ to produce a global solution of $x$ since $x_{i}$ could have multiple replicates if it is involved in several constraints. As a result, the weight, i.e., $w_{i}$ in Eq. (4) embodies the priority of a constraint – a bigger $w_{i}$ (relative to other constraints) means the global solve favors $x$ being closer to the corresponding $y_{i}$. In an extreme case when $w_{i}\rightarrow\infty$, $x_{i}\rightarrow y_{i}$ ignoring all the other constraints. One should not confuse $w_{i}$ with the stiffness of the constraint: a high constraint stiffness produces a big internal force, which could overshoot and must be coupled with a line search for extra safeguards.
@@
Here, $\hat{d}$ is a user-provided tolerance of the collision resolution. $d_{i}(x)$ denotes the closest distance between the $i$-th pair of surface primitives, either a vertex-triangle pair or an edge-edge pair. $B(d_{i})$ diverges if $d_{i}<\hat{d}$ and approaches $\infty$ when $d_{i}\rightarrow 0$. Consequently, as long as we keep $d_{i}$ positive at the beginning of a time step (i.e., all primitives are separate), the existence of $B(d_{i})$ prevents any future inter-penetration with an increasingly stronger repulsion. Lan et al. (2022b) further showed that IPC barrier function can also be integrated into LG iterations by setting the weighting function $w_{i}$ of the collision constraint as $B(d_{i})$.
@@ 4. Non-distance Barrier
While IPC (Li et al., 2021a) and its PD variations (Lan et al., 2022b) offer robust treatment for collisions, they are all based on $d_{i}$. Here, we name this family of barrier functions as the distance-based barrier or DBB. Updating DBB is expensive – one needs to perform a broad phrase collision culling to generate the list of vertex-triangle or edge-edge pairs, compute TOI for each pair, identify the smallest TOI which is between zero and one of all pairs, compute $d_{i}$, and eventually obtain the latest value of $\sum B(d_{i})$. Such a full CCD procedure is invoked frequently during LG iterations (Lan et al., 2022b) and becomes the dominant computation along the pipeline.
@@
We argue that not all DBBs are indispensable, and most of them can be replaced with a more economic alternative model under the PD framework. The underlying philosophy of DBB is to offer a nonlinear penalty, which becomes stiffer when the collision constraint is about to be violated. DBB geometrically approximates the indicator functions ($\delta_{\mathcal{A}}(x)=0$, if $x\in\mathcal{A}$; $\delta_{\mathcal{A}}(x)=\infty$, if $x\notin\mathcal{A}$). A DBB itself is not physically accurate i.e., the gradient of DBB differs from the actual collision force unless $\kappa$ in Eq. (6) is close to zero (i.e., the complementary slackness is sufficiently satisfied). In practice, what we want is an increasingly strong repulsion to correct the constraint violation, and this goal can be enabled without $d_{i}$, the actual distance between a pair of primitives. To this end, we design a non-distance barrier weight with exponential formulation:
(7) $w_{i}=B_{i}^{NDB}=k{K^{a_{i}}},$
@@
where $a_{i}\in\mathbb{Z}^{+}$ represents how many consecutive LG iterations has the constraint $C_{i}$ been active. A contact constraint is considered active if the primitive pair remains sufficiently close to each other. $k$ is an initial weight, and $K$ is the base of the growth rate. We use NDB to denote this simplified barrier model. Similar to DBB, NDB approaches to $\infty$ if a collision constraint $C_{i}$ remains unresolved over several LG iterations. In this case, the local target position of the constraint shall be satisfied with the highest priority in the global solve. Unlike DBB on the other hand, NDB no longer depends on the distance between primitives or any physically/geometrically meaningful measures. It becomes a self-adjusting variable as the optimization proceeds with minimum computation costs. As we elaborate in the next subsection, this feature delivers substantial convenience and speedup for cloth simulation.
Discussion
DBB takes the primitive distance $d(x)$ as its parameter. In contrast, our non-distance barrier (NDB) strategy does not explicitly depend on the collision distance. Instead, it employs the collision’s life span (i.e., number of iterations) to adaptively increase the weight of the constraint. NDB is not a barrier function in a strict sense since NDB is not explicitly related to $x$. Nevertheless, it introduces a penalty that asymptotically tends to infinity to prevent the interpenetration between primitives. This feature ensures the optimization variable $x$ remains within the feasible region, and works in a similar way as the conventional interior-point method. One limitation, however, is the accuracy of such a weighting mechanism. Since the iteration index is a discrete count, it is sometimes challenging to produce sufficiently different repulsion forces to distinguish nearby vertices. More iterations are therefore needed.
@@ 4.1. Partial CCD
An immediate advantage of using NDB is the alleviation of computational effort required for CCD. When updating $B_{i}^{NDB}$, it is sufficient to determine whether a collision remains active after the previous LG iteration, requiring a true-or-false response rather than an exact time of impact. This characteristic enables a more streamlined approach for NDB CCD, transforming the cubic root-finding problem into a series of dot product calculations, a technique we refer to as partial CCD.
@@
As the cloth is less stretchable, we assume that edge lengths of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ are bounded by $L$. Given a finite velocity, the distance between $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ is bounded above by $H^{1}$ during $t\in(0,1]$, and bounded below by $H^{0}$ at $t=0$. We can then obtain an upper bound of the norm of $\nabla_{\lambda}Q$:
@@
When $\rho_{\mathcal{S}}$ is smaller than $\rho^{*}$, there is at least one sample point sitting in a neighborhood of $Q<0$. Here, $\alpha<1$ is the hyperparameter used in line search filtering, which is typically set as $0.8$.
@@ Discussion
Partial CCD is sample-based and can be considered as an approximation of exact CCD algorithms (Brochu et al., 2012). We would like to mention that our method does not exclusively rely on partial CCD. In other words, partial CCD, together with exponential-based NDB collision constraint projection works as a warm start, and regular CCDs are still performed during the simulation (but at a much lower frequency). Therefore, we do not need to set the sampling density to $\rho^{*}$ in practice (as this theoretical bound depends on various factors and varies at each iteration). The performance and the quality of the simulation are not sensitive to partial CCD accuracy (e.g., see Fig. 9).
Eq. (15) appears sensitive to $H^{0}$, the closest distance between $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ for $t\in(0,1]$. For instance, if a vertex is in the proximity of the triangle at $t=0$ and hits the triangle at $t=t^{*}$, $\rho(\lambda^{*})$ of the corresponding contact approaches zero i.e., the disc of $\Omega(\lambda^{*})$ shrinks to a point. This issue can be easily fixed by setting the projection of the vertex on the triangle as a sample point if they are close to each other. In some extreme cases, where the cloth is substantially stretched or accelerated, we increase the sample density by scaling the sample interval by the change ratio of $\rho^{*}$.
A quick benchmark well demonstrates the potential of such a sample-based strategy: solving a cubic equation on the GPU as in the original IPC implementation (Li et al., 2020) is slower than evaluating Eq. (9) on a RTX 3090 GPU by over $6,000$ times. Partial CCD is about $2,000$ times faster than the state-of-the-art polynomial solver (Yuksel, 2022). For a high-resolution cloth model, we do not need a lot of sample points on a primitive, making partial CCD faster than regular CCD by orders of magnitude.
@@ 5. Subspace Reuse
NDB enables a more affordable collision processing, but it does not alter the fact that the optimization of Eq. (5) is highly nonlinear. Solving the system within a limited time budget remains a computational challenge for time-sensitive applications. Under the framework of PD, the bottleneck is the linear solve at the global stage i.e., Eq. (4). Commonly used strategies resort to iterative GPU solvers such as Jacobi (Wang, 2015) or Gauss-Seidel (Fratarcangeli et al., 2016) to solve Eq. (4) inexactly. Unfortunately, the presence of barriers makes those solvers less beneficial. Specifically, it is known that iterative solvers are effective in smoothing high-frequency errors but become cumbersome when dealing with low-frequency residuals. In contrast, low-frequency cloth deformation can be efficiently handled using subspace methods. A concrete example is shown in Figs. 2 and 3. To this end, we tackle Eq. (4) using the reduced direct solver and fullspace iterative solver, aiming to reap the goods from both sides.
@@
(16) $\mathsf{U}^{\top}\mathsf{H}\left(X+\mathsf{U}q\right)=\mathsf{U}^{\top}b$ $\;\Rightarrow\;\left(\mathsf{U}^{\top}\mathsf{H}\mathsf{U}\right)q=\mathsf{U}^{\top}\left(b-\mathsf{H}X\right)$
$\;\Rightarrow\;\Lambda q=\mathsf{U}^{\top}\left(b-\mathsf{H}X\right),$
@@
Figure 2. Bending strips w. and w/o subspace.  We simulate a collision-free scene where a cloth strip bends under gravity. The model consists of $30$K DOFs. The resulting deformation is of low frequency, which is challenging for Jacobi-like methods. A subspace solve effectively resolves this issue: only $20$ A-Jacobi iterations are needed to fully converge the simulation, which otherwise takes over one thousand iterations. Because the bending stiffness is quite strong in this example, we need to assign a big SOR-like weight ($\omega=0.9$) to dampen each A-Jacobi iteration. Our method runs over $300$ FPS for this example, while PD-IPC (Lan et al., 2022b) is less than 0.5 FPS due to the large number of A-Jacobi iterations.
@@
Figure 3. Spectral distribution of residual errors (w/o collision).  We plot the distribution of residual error over the first 100 modal bases of the strip test (Fig. 2). As shown at the top, the dominant deformation is low-frequency, which is efficiently solved within the subspace (bottom). On the other hand, A-Jacobi iterations are not effective in dealing with low-frequency residuals. 100 A-Jacobi iterations barely lower the low-frequency errors, while the high-frequency strains are well relaxed (middle).
@@
As $\mathsf{U}$ is eigenvectors, $\mathsf{U}^{\top}\mathsf{H}\mathsf{U}=\Lambda$ becomes a diagonal matrix of eigenvalues. If the constraint set stays unchanged during the simulation, solving Eq. (16) is highly efficient on the GPU – we only need one subspace projection for evaluating $\mathsf{U}^{\top}(b-\mathsf{H}X)$, and diagonalized subspace solve is negligible. The resulting subspace displacement of $q$ is then converted to its fullspace counterpart as $u=\mathsf{U}q$. At this point, most low-frequency errors have been eliminated by the subspace solve, and $\widetilde{x}=X+\mathsf{U}q$ represents an ideal guess of Eq (4) for iterative solvers, which only has high-frequency errors. We then use aggregated-Jacobi (A-Jacobi) as in (Lan et al., 2022b) to solve for $\Delta x$ to relax the (high-frequency) residual of $\widetilde{x}$:
(17) $\mathsf{H}(\widetilde{x}+\Delta x)=b\;\Rightarrow\;\mathsf{H}\Delta x=b-\mathsf{H}\widetilde{x}\;\Rightarrow\;\mathsf{H}\Delta x=b-\mathsf{H}(X+\mathsf{U}q).$
@@
A representative experiment is reported in Fig. 2, where a cloth strip is attached to the wall and gets bent under gravity. When the low-frequency deformation is computed within the subspace ($r=30$), it only takes 20 A-Jacobi iterations to converge the simulation to the ground truth i.e., the exact global solve. General A-Jacobi iterations are not effective for such low-frequency deformations – we observe noticeable visual difference even after 600 iterations (the right-most beam). In this case, the bending stiffness of the cloth is relatively high, and the vanilla Jacobi or A-Jacobi do not even converge. We use a large successive over-relaxation-like (SOR) weight ($\omega=0.9$) to dampen each A-Jacobi update. The error distribution of this experiment is plotted in Fig. 3, which is consistent with our previous analysis.
Figure 4. Subspace reuse.   A piece of tablecloth (66K DOFs) drops on a wooden Armadillo. Over $30\%$ of vertices are involved in collision constraints.
Collisions are ubiquitous in cloth simulation, and $\mathsf{H}$ varies during the time integration. Building a new subspace at the simulation runtime does not sound practical for interactive simulations. Our key observation here is: while the global-stage matrix is altered by different collision constraints, the essential structure of its low-frequency subspace remains unchanged. It seems counterintuitive at first sight: how can old eigenvectors still be effective while the matrix is modified? This is because low-frequency modes built at the rest shape depict general and global strain distribution over the garment. This is not strongly coupled with high-frequency deformations. For example, the appearance of local wrinkles is less influenced by the overarching movements of the cloth but more by the specific configurations of the corresponding collisions.
Interestingly, it is the low-frequency deformations that most severely hinder the convergence of iterative solvers. By eliminating or even just reducing these low-frequency errors, we could significantly enhance the convergence of the subsequent Jacobi method.
@@
Figure 6. Convergence curves w. and w/o subspace (reuse).  We plot the convergence curves for experiments in Figs. 2 and 4. We can see subspace solve saves a large faction of A-Jacobi iterations even under intensive collisions. The computation cost of one subspace solve, on the other hand, is similar to performing one or two A-Jacobi iterations on average.